OpenStudy (anonymous):

evaluate the integral: xe^(2x^2) dx

OpenStudy (anonymous):

Let u=2x^2 then du=4xdx so (1/4)du=xdx So your integral becomes: \[\frac{1}{4} \int\limits e^u du=\frac{1}{4}e^u+C=\frac{1}{4}e^{2x^2}+C\]